16. Solve the equation using the Zero Product Property.
(2x – 4)(2x – 1) = 0 (1 point)

Answer by checkley79(3341)

(2x – 4)(2x – 1) = 0
2X-4=0
2X=4
X=4/2
X=4/2
X=2 ANS.
2X-1=0
2X=1
X=1/2 OR .5 ANS,